# Unlocking Textual and Visual Wisdom: Open-Vocabulary 3D Object Detection Enhanced by Comprehensive Guidance from Text and Image

### v1

@@ 5 Conclusion
In this study, we delve into the challenging realm of open-vocabulary 3D object detection. While vision language foundation models have propelled 2D detection forward in open-vocabulary scenarios, their potential for 3D detection remains underutilized. To address this gap, we introduce a novel framework called Image-Guided Novel Object Discovery and Hierarchical Multi-Modality Feature Space Alignment framework. Specifically, we integrate 2D detection model and 3D detector to identify novel objects. Additionally, we hierarchically align the 3D features with the vision-language feature space at the instance, category, and scene levels. Our approach capitalizes the power of foundation models to extract comprehensive guidance information from both text and images, which is then effectively integrated with 3D inputs for open-vocabulary 3D object detection. Extensive experiments validate the effectiveness of our proposed methods, demonstrating the efficacy in various datasets.
Acknowledgments This research work is supported by the Agency for Science, Technology and Research (A*STAR) under its MTC Programmatic Funds (Grant No. M23L7b0021).

### v2

@@ 5 Conclusion
In this study, we delve into the challenging realm of open-vocabulary 3D object detection. While vision language foundation models have propelled 2D detection forward in open-vocabulary scenarios, their potential for 3D detection remains underutilized. To address this gap, we introduce a novel framework: the image-guided novel class discovery and hierarchical feature space alignment framework, dubbed as INHA. Specifically, we integrate 2D detection model and 3D detector to discover novel objects. Additionally, we hierarchically align the 3D features with the vision-language feature space at the instance, category, and scene levels. Our INHA capitalizes the power of foundation models to extract comprehensive guidance information from both text and images, which is then effectively integrated with 3D inputs for open-vocabulary 3D object detection. Extensive experiments validate the effectiveness of our INHA across various datasets.
Acknowledgments This research work is partially supported by the Shanghai Science and Technology Program (Project No. 21JC1400600), and the Agency for Science, Technology and Research (A*STAR) under its MTC Programmatic Funds (Grant No. M23L7b0021).